Sorry I forgot to write out the pair of lines
3x+8=y
2y=6x-3
are these lines parallel?

Found 2 solutions by mananth, Maths68:
Answer by mananth(16946)

3x+8=y
y=3x+8

2y=6x-3
/2
y=3x-(3/2)
compare both the lines with y= mx+b
m the slope of both lines are 3 so the lines are parallel

Answer by Maths68(1474)

3x+8=y............(1)
y=3x+8
Compare above equation with the standard equation of the line
y=mx+b
slope =m=3 and y-intercept =b=8
2y=6x-3...........(2)
Divide above equation by 2
2y/2=(6x-3)/2
y=6x/2-3/2
y=3x-3/2
Compare above equation with the standard equation of the line
y=mx+b
slope =m=3 and y-intercept =b=-3/2


Both lines are parallel because their slopes are same.